After plotting demand for four periods, an emergency room manager has concluded that a trend-adjusted exponential smoothing model is appropriate to predict future demand. The initial estimate of trend is based on the net change of 27 for the three periods from 1 to 4, for an average of +9.00 units.

Period Actual Period Actual
1 206 6 263
2 233 7 278
3 228 8 288
4 233 9 293
5 253 10


Use α=.50 and β=.10, and TAF of 245 for period 5. Obtain forecasts for periods 6 through 10. (Round your intermediate calculations and final answers to 2 decimal places.)

t Period TAFt
6
7
8
9
10
Answer #1

Answers are in BOLD

Year Period, X Number Exponential smoothing, Ft Trend, Tt forecast including trend, FITt (TAFt)
2009 1 206
2010 2 233
2011 3 228
2012 4 233
2013 5 253 9.00 245.00
2014 6 263 249.00 9.40 258.40
2015 7 278 260.70 9.63 270.33
2016 8 288 274.17 10.01 284.18
2017 9 293 286.09 10.20 296.29
2018 10 294.65 10.04 304.69
  • Alpha= 0.5, Beta= 0.1
  • the formula to be used in Exponential smoothing, Ft is

Ft+1= FITt + alpha*(At+1 – FITt)

FITt is Forecast including trend, Ft+1 means Exponential forecast demand of t+1'th period, if you want to find out the Forecast of 3rd period = Forecast including trend of 2nd period + alpha*(Actual demand of 3rd period – Forecast including trend of 2nd period)

remember FIT5 = 245

  • the formula to be used in Trend smoothing, Tt is

Tt+1= Tt + beta*(Ft+1 – FITt)

Ft+1 means Exponential forecast demand of t+1'th period, if you want to find out the Trend smoothing Forecast = forecast of 3rd period = Trend smoothing Forecast of 2nd period + beta*(Exponential forecast of 3rd period – Forecast including trend of 2nd period)

remember Trend of 5th period is (Actual of 4- Actual of 1)/3= (233-206)/3= 9

  • the formula to be used in Forecast including Trend, FIT is

FITt= Tt + Ft